The length of a rectangle is 7 meters longer than its width. What is the width of this rectangle if its perimeter is equal to 150 meters?
Answer by checkley77(12844)

L=W+7
2L+2W=P
2(W+7)+2W=150
2W+14+2W=150
4W=150-14
4W=136
W=136/4
W=34 METERS IS THE LENGTH.
L=34+7=41 METERS IS THE LENGTH.
PROOF:
2*41+2*34=150
82+68=150
150=150